OpenStudy (surana):

Find the range of the function f(x) /- 12x+24

OpenStudy (surana):

Technically, it's \[f(x)\sqrt{12x+24}\]

OpenStudy (surana):

Possible answers are... A: \[y \le 12\] B: \[y \ge 24\] C: \[y \le 0 \] D: \[y \ge 0\]

OpenStudy (jdoe0001):

hmm have you covered parabolas yet?

OpenStudy (surana):

This is radicals, but I've covered parabolas to some extent.

OpenStudy (surana):

Why?

OpenStudy (jdoe0001):

one sec

OpenStudy (jdoe0001):

ok.... so.... the range the range means, the "y values", or the values, "y" does take, based on whatever "x" is so \(\bf f(x)=y=\sqrt{12x+24}\implies y^2=12x+24 \\ \quad \\ y^2-0x-0^2=12x+24\implies (y-0)^2=12x+24 \\ \quad \\ (y-0)^2=12(x+2)\impliedby \textit{"y" is squared, meaning a horizontal parabola}\) any ideas on what the vertex for that parabola is? where does it open towards? left or right?

OpenStudy (surana):

I think it's supposed to open to the left?

OpenStudy (jdoe0001):

the parabola is horizontal has a vertex, or U-turn "p" is positive, that is , the common factor for the "x", or 12, is positive, that means, opens to the right

OpenStudy (surana):

Ah, so it opens to the right then, so what happens next?

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

the "domain" is restricted, since it's inside a root so, it cannot make the root value negative, so "x" cannot be, as you can see in the graph less than -2, otherwise the radicand turns negative, and the root becomes imaginary but that's the domain

OpenStudy (surana):

So if it's from positive infinity to negative infinity, it could be C or D?

OpenStudy (jdoe0001):

the range... if you notice, just keeps on upwards and downwards

OpenStudy (surana):

So it could be C? Equal to or less than zero?

OpenStudy (jdoe0001):

well.... I think none of those choices are.... fit since any of them are correct one can say, since the "range" as opposed to the domain has a set of \(\large y \in \mathbb{R}\)

OpenStudy (surana):

So which one would be the best choice?

OpenStudy (jdoe0001):

hmm

OpenStudy (surana):

Ah nuts, I just realized the thing was written the way it was, but with a = between fx and the rest of the problem. So it's the equation I wrote above, but with an = between Fx and the rest of it. Sorry.

OpenStudy (surana):

Hopefully it won't impact it too much.

OpenStudy (jdoe0001):

same thing... same equation really, yes, I noticed that \(\bf f(x)=y=\sqrt{12x+24}\implies y^2=12x+24 \\ \quad \\ y^2-0x-0^2=12x+24\implies (y-0)^2=12x+24 \\ \quad \\ (y-0)^2=12(x+2)\impliedby \textit{"y" is squared, meaning a horizontal parabola}\)

OpenStudy (surana):

Oh good. Thank you. I was worried for a minute that you were going to get mad at me for wasting your time. So it's still the same thing?

OpenStudy (jdoe0001):

nahh.... but.... hmmm the range is that "y" goes from \(\large (+\infty, -\infty)\)

OpenStudy (surana):

So that means it could be one or the other?

OpenStudy (jdoe0001):

hmmm lemme decompose it further.... I think it has a constraint

OpenStudy (jdoe0001):

\(\bf f(x)=y=\sqrt{12x+24}\implies y^2=12x+24 \\ \quad \\ y^2-24=12x\implies \cfrac{y^2-24}{12}=x\) notice that based on that, I don't see any constraints on "y" per se it could be negative anything, or positive anything the square exponential, will make any negaative positive, or any positive positive anyway

OpenStudy (surana):

So it could be pretty much anything?

OpenStudy (jdoe0001):

yeap... notice the graph, the parabola just keeps on going upwards and downwards slowly but surely the one with the constraints is mainly the "domain", or the "x" value

OpenStudy (surana):

So what's your opinion on what would be the best choice?

OpenStudy (jdoe0001):

you could try reposting .... or tag someone else... maybe they have a different take on it but I'm looking at the graph, and the range is just infinity to infinity is it y>0 or y<12 or y>24 or whatever else, yes yes and yes

OpenStudy (surana):

So I'm stuck one way or another?

OpenStudy (surana):

I'll go with the 0 one then.

OpenStudy (jdoe0001):

hhehe, sure, but basically, any choice will do

OpenStudy (surana):

Yeah, I suppose if you state the answer goes off into infinity, might as well go with something that kind of comes close to it.

OpenStudy (surana):

Thanks.

OpenStudy (jdoe0001):

yw